OpenStudy (anonymous):

If f(x) = x^2/x + 3 find f(x + h).

OpenStudy (anonymous):

\[\frac{ x^2+h^2+2hx }{ x+3+h }\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[\frac{ x^2+h }{ x+3+h }\]

OpenStudy (zale101):

If \( f(x) =\Large \frac{x^2}{x + 3}\) \(f(x+h)=\Large\frac{(x+h)^2}{(x+h)+3}\)

OpenStudy (anonymous):

so its the second one I put?

OpenStudy (zale101):

\(f(x)=\Large \frac{x^2+2hx+h^2}{x+h+3}\) Yes, correct :)

OpenStudy (anonymous):

that looks like the first one^

OpenStudy (zale101):

The first one is correct